# The theory of implicit operations

### v1

@@ 13. Absolutely closed and primal algebras
To prove the inclusion $\mathsf{K}_{\textsc{ac}}\subseteq\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$, consider $\bm{A}\in\mathsf{K}_{\textsc{ac}}$. Since $\mathcal{F}\subseteq\mathsf{ext}(\mathsf{K})$, 9.6 implies that there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. We show that $A$ is the universe of a subalgebra of $\bm{B}$. From $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$ it follows that $A$ is closed under the operations of the language of $\mathsf{K}$. Recall that every operation symbol of ${L}_{\mathcal{F}}-{L}_{\mathsf{K}}$ is of the form $g_{f}$ for some $f\in\mathcal{F}$. Consider an $n$-ary $f\in\mathcal{F}$. We will show that $A$ is closed under $g_{f}^{\bm{B}}$. Since $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$, we have $g_{f}^{\bm{B}}=f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}$. Let $a_{1},\dots,a_{n}\in A$. Since $f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ is total, 4.10 yields $f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$. As $\bm{A}$ is absolutely closed in $\mathsf{K}$ by assumption, we have $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$, and hence $f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})\in A$. We have shown that $A$ is the universe of a subalgebra of $\bm{B}$. Therefore, we can expand $\bm{A}$ to an ${L}_{\mathsf{M}}$-algebra $\bm{A}^{*}$ by setting $g_{f}^{\bm{A}^{*}}=g_{f}^{\bm{B}}{\upharpoonright}_{A}$ for every $g_{f}\in{L}_{\mathcal{F}}-{L}_{\mathsf{K}}$. The definition of $\bm{A}^{*}$ guarantees that $\bm{A}^{*}\leqslant\bm{B}$. Since $\bm{B}\in\mathsf{M}$ and $\mathsf{M}$ is a universal class by 10.3(i), we obtain $\bm{A}^{*}\in\mathbb{S}(\mathsf{M})\subseteq\mathsf{M}$. Thus, we conclude that $\bm{A}=\bm{A}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$.
@@
It remains to show that, when $\mathsf{M}$ is an equational Beth companion of $\mathsf{K}$, we have $\mathsf{K}_{\textsc{ac}}=\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Suppose that $\mathsf{M}$ is an equational Beth companion of $\mathsf{K}$. Then $\mathsf{M}$ is faithfully term equivalent to a Beth companion $\mathsf{M}^{*}$ of $\mathsf{K}$ induced by a family of operations defined by conjunctions of equations. Moreover, 10.4 yields that $\mathsf{M}^{*}$ is of the form $\mathsf{K}[{L}_{\mathcal{F}^{*}}]$ for some $\mathcal{F}^{*}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$. Then $\mathsf{K}[{L}_{\mathcal{F}^{*}}]{\upharpoonright}_{{L}_{\mathsf{K}}}=\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Since $\mathsf{K}[{L}_{\mathcal{F}^{*}}]{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{K}_{\textsc{ac}}\subseteq\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$ by the first two paragraphs of this proof, we have $\mathsf{K}_{\textsc{ac}}=\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\mathsf{M}$ and $\mathsf{M}^{*}$ are Beth companions of $\mathsf{K}$, 13.29 yields $\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}=\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Hence, we conclude that $\mathsf{K}_{\textsc{ac}}=\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$. ∎
Lastly, we prove 13.3.
@@
In view of 13.28(ii), to prove that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, it suffices to show that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$ for all $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. To this end, consider $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. Since $\mathsf{M}$ is a universal class by 10.3(i), we obtain $\bm{A}\in\mathsf{M}$. Therefore, the assumption that $\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{K}_{\textsc{ac}}$ implies $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{ac}}$. Hence, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. ∎
@@ 14. Classes without a Beth companion
We close this work by providing some examples of classes of algebras lacking a Beth companion. The two main results of the section concern the varieties of monoids and of commutative monoids (14.1), and certain quasivarieties of Heyting algebras (14.11). We begin with the result on monoids.
@@
In order to prove 14.1, we first need to introduce the notion of a dominion base and establish some technical results about dominion bases and implicit operations.
@@
Let $\mathsf{K}$ be a class of algebras and $\Delta\subseteq\mathsf{imp}_{pp}(\mathsf{K})$. We say that $\Delta$ is a dominion base for $\mathsf{K}$ when for all $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$ there exist $f\in\Delta$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}({f^{\bm{B}}})\cap A^{n}$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=b$.
4.10 states that $\mathsf{imp}_{pp}(\mathsf{K})$ is a dominion base for every elementary class $\mathsf{K}$, and 4.9 states that Isbell’s operations (see 3.14) form a dominion base for the varieties of monoids and of commutative monoids. The following result illustrates how having a concrete and transparent dominion base simplifies the task of finding interpolants for implicit operations.
@@
Let $\mathsf{K}$ be a quasivariety with dominion base $\Delta$ and $f\in\mathsf{imp}_{pp}(\mathsf{K})$ of arity $n$. Then there exist $g\in\Delta$ and $n$-ary terms $t_{1},\dots,t_{m}$ of $\mathsf{K}$ such that the composition $g(t_{1}^{\mathsf{K}},\dots,t_{m}^{\mathsf{K}})$ interpolates $f$ in $\mathsf{K}$.
@@
where $\psi$ is a conjunction of equations. Let $X=\{z_{1},\dots,z_{k},x_{1},\dots,x_{n},y\}$. Since $\mathsf{K}$ is a quasivariety, the free algebra $\bm{F}_{\mathsf{K}}(X)$ belongs to $\mathsf{K}$ (see 2.19). We denote by $\theta$ the $\mathsf{K}$-congruence of $\bm{F}_{\mathsf{K}}(X)$ generated by the pairs $\langle s_{1},s_{2}\rangle$, where $s_{1}\thickapprox s_{2}$ is an equation in $\psi$. Consider $\bm{B}=\bm{F}_{\mathsf{K}}(X)/\theta$ and $\bm{A}=\mathsf{Sg}^{\bm{B}}(x_{1}/\theta,\dots,x_{n}/\theta)$. The definition of $\theta$ implies that
@@
Therefore, $\langle x_{1}/\theta,\dots,x_{n}/\theta\rangle\in\mathsf{dom}(f^{\bm{B}})$ and $f^{\bm{B}}(x_{1}/\theta,\dots,x_{n}/\theta)=y/\theta$ because $f$ is defined by $\varphi$. As $f\in\mathsf{imp}_{pp}(\mathsf{K})$ by assumption and $x_{1}/\theta,\dots,x_{n}/\theta\in A$ by the definition of $\bm{A}$, it follows from 4.10 that $y/\theta\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. Since $\Delta$ is a dominion base for $\mathsf{K}$ and $\bm{A}$ is generated by $x_{1}/\theta,\dots,x_{n}/\theta$, there exist an $m$-ary $g\in\Delta$ and terms $t_{1}(x_{1},\dots,x_{n}),\dots,t_{m}(x_{1},\dots,x_{n})$ of $\mathsf{K}$ such that
@@
for every equation $s_{1}\thickapprox s_{2}$ in $\psi$. As $X=\{z_{1},\dots,z_{k},x_{1},\dots,x_{n},y\}$ is a set of free generators for $\bm{F}_{\mathsf{K}}(X)$, there exists a homomorphism $h\colon\bm{F}_{\mathsf{K}}(X)\to\bm{C}$ such that $h(z_{i})=e_{i}$ and $h(x_{i})=c_{i}$ for each $i$, and $h(y)=d$. The definition of $\theta$ and (66) yield $\theta\subseteq\mathsf{Ker}(h)$. Since $\bm{B}=\bm{F}_{\mathsf{K}}(X)/\theta$, 2.6 implies that the homomorphism $h$ induces a homomorphism $k\colon\bm{B}\to\bm{C}$ such that $k(z_{i}/\theta)=e_{i}$ and $k(x_{i}/\theta)=c_{i}$ for each $i$, and $k(y/\theta)=d$. Since $k$ is a homomorphism, we obtain
@@
for each $i\leqslant m$. So, (64) implies $\langle t_{1}^{\bm{C}}(c_{1},\dots,c_{n}),\dots,t_{m}^{\bm{C}}(c_{1},\dots,c_{n})\rangle\in\mathsf{dom}(g^{\bm{C}})$ because $g\in\mathsf{imp}_{pp}(\mathsf{K})$. We also have
@@
where the first equality holds because $k(x_{i}/\theta)=c_{i}$ for each $i$, the second follows from the assumptions that $k$ is a homomorphism and $g\in\mathsf{imp}_{pp}(\mathsf{K})$, the third from (65), and the last from $k(y/\theta)=d$. Hence, we conclude that $g^{\bm{C}}(t_{1}^{\bm{C}}(c_{1},\dots,c_{n}),\dots,t_{m}^{\bm{C}}(c_{1},\dots,c_{n}))=d$, as desired. ∎
We now apply 14.3 and the fact that Isbell’s operations form a dominion base to deduce a useful property of implicit operations of the variety of (commutative) monoids.
@@
Let $\mathsf{K}$ be either the variety of monoids or the variety of commutative monoids. For every unary $f\in\mathsf{imp}_{pp}(\mathsf{K})$ there exist $l,r\in\mathbb{N}$ such that $a^{l}f^{\bm{A}}(a)=a^{r}$ for all $\bm{A}\in\mathsf{K}$ and $a\in\mathsf{dom}(f^{\bm{A}})$.
@@
Consider a unary $f\in\mathsf{imp}_{pp}(\mathsf{K})$. For each $n\in\mathbb{N}$ we denote by $g_{n}$ the implicit operation of $\mathsf{K}$ defined by the $n$-th Isbell’s formula (see 3.14). As Isbell’s operations form a dominion base for $\mathsf{K}$, 14.3 yields $n\in\mathbb{N}$ and unary terms $t_{1},\dots,t_{2n+1}$ such that $g_{n}(t_{1}^{\mathsf{K}},\dots,t_{2n+1}^{\mathsf{K}})$ interpolates $f$ in $\mathsf{K}$. As unary terms of $\mathsf{K}$ are equivalent to powers of a variable, there exist $e_{1},\dots,e_{n+1}\in\mathbb{N}$ such that $\mathsf{K}\vDash t_{i}(x)\thickapprox x^{e_{i}}$ for each $i\leqslant 2n+1$.
@@
Let $\mathsf{K}$ be a quasivariety with a Beth companion of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Then for every $f\in\mathsf{imp}_{pp}(\mathsf{K})$ there exists $g\in\mathsf{ext}_{pp}(\mathsf{K})$ that interpolates $f$ in $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$.
@@
Since $f\in\mathsf{imp}_{pp}(\mathsf{K})$, by 11.6 there exists a term $t$ of ${L}_{\mathcal{F}}$ that interpolates $f$ in $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As $f$ has positive arity and $t$ interpolates $f$, it follows that $t$ is not a constant. So, 10.22(ii) yields $g\in\mathsf{ext}_{pp}(\mathsf{K})$ such that
@@
Let $\mathsf{K}$ be either the variety of monoids or the variety of commutative monoids and assume, with a view to contradiction, that $\mathsf{K}$ has a Beth companion. 11.5 implies that $\mathsf{K}$ has also a Beth companion of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Let $f$ be the operation of taking inverses in monoids. Then $f\in\mathsf{imp}_{pp}(\mathsf{K})$ by 3.7. 14.6 yields a unary $g\in\mathsf{ext}_{pp}(\mathsf{K})$ that interpolates $f$ in $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$. By 14.4 there exist $l,r\in\mathbb{N}$ such that $a^{l}g^{\bm{A}}(a)=a^{r}$ for all $\bm{A}\in\mathsf{K}$ and $a\in\mathsf{dom}(g^{\bm{A}})$. Since $\mathsf{K}$ is a variety and $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ a pp expansion of $\mathsf{K}$, 10.2 implies that $\mathsf{K}=\mathbb{S}((\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])){\upharpoonright}_{{L}_{\mathsf{K}}})$. Hence, $\mathsf{K}\subseteq\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}})$. Therefore, there exists an extension $\bm{B}$ of the multiplicative monoid $\mathbb{Q}$ of the rationals such that $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$. Then $g^{\bm{B}}$ is total because $g\in\mathsf{ext}_{pp}(\mathsf{K})=\mathcal{F}$. Since $g$ interpolates $f$ in $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$, we obtain $g^{\bm{B}}(2)=f^{\bm{B}}(2)=f^{\mathbb{Q}}(2)=2^{-1}$. Therefore, $2^{l-1}=2^{l}g^{\bm{B}}(2)=2^{r}$. So, $2^{l-1}=2^{r}$ holds in $\mathbb{Q}$. In turn, this implies $l=r+1$. Consequently,
@@
Let $\bm{C}$ and $c\in C$ be as in 14.7. Since $g\in\mathsf{ext}_{pp}(\mathsf{K})$, there exists $\bm{D}\in\mathsf{K}$ such that $\bm{C}\leqslant\bm{D}$ and $g^{\bm{D}}$ is total. From $c\in\mathsf{dom}(g^{\bm{D}})$ and (68) it follows $c^{r+1}g^{\bm{D}}(c)=c^{r}$. Thus, using $c^{r+1}=c^{r+2}$, we deduce

### v2

@@ 13. Absolutely closed and primal algebras
To prove the inclusion $\mathsf{K}_{\textsc{ac}}\subseteq\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$, consider $\bm{A}\in\mathsf{K}_{\textsc{ac}}$. Since $\mathcal{F}\subseteq\mathsf{ext}(\mathsf{K})$, Proposition 9.6 implies that there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. We show that $A$ is the universe of a subalgebra of $\bm{B}$. From $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$ it follows that $A$ is closed under the operations of the language of $\mathsf{K}$. Recall that every operation symbol of ${L}_{\mathcal{F}}-{L}_{\mathsf{K}}$ is of the form $g_{f}$ for some $f\in\mathcal{F}$. Consider an $n$-ary $f\in\mathcal{F}$. We will show that $A$ is closed under $g_{f}^{\bm{B}}$. Since $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$, we have $g_{f}^{\bm{B}}=f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}$. Let $a_{1},\dots,a_{n}\in A$. Since $f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ is total, Theorem 4.10 yields $f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$. As $\bm{A}$ is absolutely closed in $\mathsf{K}$ by assumption, we have $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$, and hence $f^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})\in A$. We have shown that $A$ is the universe of a subalgebra of $\bm{B}$. Therefore, we can expand $\bm{A}$ to an ${L}_{\mathsf{M}}$-algebra $\bm{A}^{*}$ by setting $g_{f}^{\bm{A}^{*}}=g_{f}^{\bm{B}}{\upharpoonright}_{A}$ for every $g_{f}\in{L}_{\mathcal{F}}-{L}_{\mathsf{K}}$. The definition of $\bm{A}^{*}$ guarantees that $\bm{A}^{*}\leqslant\bm{B}$. Since $\bm{B}\in\mathsf{M}$ and $\mathsf{M}$ is a universal class by Theorem 10.3(i), we obtain $\bm{A}^{*}\in\mathbb{S}(\mathsf{M})\subseteq\mathsf{M}$. Thus, we conclude that $\bm{A}=\bm{A}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$.
@@
It remains to show that, when $\mathsf{M}$ is an equational Beth companion of $\mathsf{K}$, we have $\mathsf{K}_{\textsc{ac}}=\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Suppose that $\mathsf{M}$ is an equational Beth companion of $\mathsf{K}$. Then $\mathsf{M}$ is faithfully term equivalent to a Beth companion $\mathsf{M}^{*}$ of $\mathsf{K}$ induced by a family of operations defined by conjunctions of equations. Moreover, Theorem 10.4 yields that $\mathsf{M}^{*}$ is of the form $\mathsf{K}[{L}_{\mathcal{F}^{*}}]$ for some $\mathcal{F}^{*}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$. Then $\mathsf{K}[{L}_{\mathcal{F}^{*}}]{\upharpoonright}_{{L}_{\mathsf{K}}}=\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Since $\mathsf{K}[{L}_{\mathcal{F}^{*}}]{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{K}_{\textsc{ac}}\subseteq\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$ by the first two paragraphs of this proof, we have $\mathsf{K}_{\textsc{ac}}=\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\mathsf{M}$ and $\mathsf{M}^{*}$ are Beth companions of $\mathsf{K}$, Proposition 13.29 yields $\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}=\mathsf{M}^{*}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Hence, we conclude that $\mathsf{K}_{\textsc{ac}}=\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}$. ∎
Lastly, we prove Theorem 13.3.
@@
In view of Proposition 13.28(ii), to prove that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, it suffices to show that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$ for all $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. To this end, consider $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. Since $\mathsf{M}$ is a universal class by Theorem 10.3(i), we obtain $\bm{A}\in\mathsf{M}$. Therefore, the assumption that $\mathsf{M}{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{K}_{\textsc{ac}}$ implies $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{ac}}$. Hence, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. ∎
@@ 14. Classes without a Beth companion
We close this work by providing some examples of classes of algebras lacking a Beth companion. The two main results of the section concern the varieties of monoids and of commutative monoids (Theorem 14.1), and certain quasivarieties of Heyting algebras (Theorem 14.11). We begin with the result on monoids.
@@
In order to prove Theorem 14.1, we first need to introduce the notion of a dominion base and establish some technical results about dominion bases and implicit operations.
@@
Let $\mathsf{K}$ be a class of algebras and $\Delta\subseteq\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. We say that $\Delta$ is a dominion base for $\mathsf{K}$ when for all $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$ there exist $f\in\Delta$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}({f^{\bm{B}}})\cap A^{n}$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=b$.
Theorem 4.10 states that $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ is a dominion base for every elementary class $\mathsf{K}$, and Isbell’s Zigzag Theorem 4.9 states that Isbell’s operations (see Example 3.14) form a dominion base for the varieties of monoids and of commutative monoids. The following result illustrates how having a concrete and transparent dominion base simplifies the task of finding interpolants for implicit operations.
@@
Let $\mathsf{K}$ be a quasivariety with dominion base $\Delta$ and $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ of arity $n$. Then there exist $g\in\Delta$ and $n$-ary terms $t_{1},\dots,t_{m}$ of $\mathsf{K}$ such that the composition $g(t_{1}^{\mathsf{K}},\dots,t_{m}^{\mathsf{K}})$ interpolates $f$ in $\mathsf{K}$.
@@
where $\psi$ is a conjunction of equations. Let $X=\{z_{1},\dots,z_{k},x_{1},\dots,x_{n},y\}$. Since $\mathsf{K}$ is a quasivariety, the free algebra $\bm{F}_{\mathsf{K}}(X)$ belongs to $\mathsf{K}$ (see Theorem 2.19). We denote by $\theta$ the $\mathsf{K}$-congruence of $\bm{F}_{\mathsf{K}}(X)$ generated by the pairs $\langle s_{1},s_{2}\rangle$, where $s_{1}\thickapprox s_{2}$ is an equation in $\psi$. Consider $\bm{B}=\bm{F}_{\mathsf{K}}(X)/\theta$ and $\bm{A}=\mathsf{Sg}^{\bm{B}}(x_{1}/\theta,\dots,x_{n}/\theta)$. The definition of $\theta$ implies that
@@
Therefore, $\langle x_{1}/\theta,\dots,x_{n}/\theta\rangle\in\mathsf{dom}(f^{\bm{B}})$ and $f^{\bm{B}}(x_{1}/\theta,\dots,x_{n}/\theta)=y/\theta$ because $f$ is defined by $\varphi$. As $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ by assumption and $x_{1}/\theta,\dots,x_{n}/\theta\in A$ by the definition of $\bm{A}$, it follows from Theorem 4.10 that $y/\theta\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. Since $\Delta$ is a dominion base for $\mathsf{K}$ and $\bm{A}$ is generated by $x_{1}/\theta,\dots,x_{n}/\theta$, there exist an $m$-ary $g\in\Delta$ and terms $t_{1}(x_{1},\dots,x_{n}),\dots,t_{m}(x_{1},\dots,x_{n})$ of $\mathsf{K}$ such that
@@
for every equation $s_{1}\thickapprox s_{2}$ in $\psi$. As $X=\{z_{1},\dots,z_{k},x_{1},\dots,x_{n},y\}$ is a set of free generators for $\bm{F}_{\mathsf{K}}(X)$, there exists a homomorphism $h\colon\bm{F}_{\mathsf{K}}(X)\to\bm{C}$ such that $h(z_{i})=e_{i}$ and $h(x_{i})=c_{i}$ for each $i$, and $h(y)=d$. The definition of $\theta$ and (66) yield $\theta\subseteq\mathsf{Ker}(h)$. Since $\bm{B}=\bm{F}_{\mathsf{K}}(X)/\theta$, Proposition 2.6 implies that the homomorphism $h$ induces a homomorphism $k\colon\bm{B}\to\bm{C}$ such that $k(z_{i}/\theta)=e_{i}$ and $k(x_{i}/\theta)=c_{i}$ for each $i$, and $k(y/\theta)=d$. Since $k$ is a homomorphism, we obtain
@@
for each $i\leqslant m$. So, (64) implies $\langle t_{1}^{\bm{C}}(c_{1},\dots,c_{n}),\dots,t_{m}^{\bm{C}}(c_{1},\dots,c_{n})\rangle\in\mathsf{dom}(g^{\bm{C}})$ because $g\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. We also have
@@
where the first equality holds because $k(x_{i}/\theta)=c_{i}$ for each $i$, the second follows from the assumptions that $k$ is a homomorphism and $g\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, the third from (65), and the last from $k(y/\theta)=d$. Hence, we conclude that $g^{\bm{C}}(t_{1}^{\bm{C}}(c_{1},\dots,c_{n}),\dots,t_{m}^{\bm{C}}(c_{1},\dots,c_{n}))=d$, as desired. ∎
We now apply Theorem 14.3 and the fact that Isbell’s operations form a dominion base to deduce a useful property of implicit operations of the variety of (commutative) monoids.
@@
Let $\mathsf{K}$ be either the variety of monoids or the variety of commutative monoids. For every unary $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ there exist $l,r\in\mathbb{N}$ such that $a^{l}f^{\bm{A}}(a)=a^{r}$ for all $\bm{A}\in\mathsf{K}$ and $a\in\mathsf{dom}(f^{\bm{A}})$.
@@
Consider a unary $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. For each $n\in\mathbb{N}$ we denote by $g_{n}$ the implicit operation of $\mathsf{K}$ defined by the $n$-th Isbell’s formula (see Example 3.14). As Isbell’s operations form a dominion base for $\mathsf{K}$, Theorem 14.3 yields $n\in\mathbb{N}$ and unary terms $t_{1},\dots,t_{2n+1}$ such that $g_{n}(t_{1}^{\mathsf{K}},\dots,t_{2n+1}^{\mathsf{K}})$ interpolates $f$ in $\mathsf{K}$. As unary terms of $\mathsf{K}$ are equivalent to powers of a variable, there exist $e_{1},\dots,e_{n+1}\in\mathbb{N}$ such that $\mathsf{K}\vDash t_{i}(x)\thickapprox x^{e_{i}}$ for each $i\leqslant 2n+1$.
@@
Let $\mathsf{K}$ be a quasivariety with a Beth companion of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Then for every $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ there exists $g\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ that interpolates $f$ in $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$.
@@
Since $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, by Theorem 11.6 there exists a term $t$ of ${L}_{\mathcal{F}}$ that interpolates $f$ in $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As $f$ has positive arity and $t$ interpolates $f$, it follows that $t$ is not a constant. So, Proposition 10.22(ii) yields $g\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ such that
@@
Let $\mathsf{K}$ be either the variety of monoids or the variety of commutative monoids and assume, with a view to contradiction, that $\mathsf{K}$ has a Beth companion. Corollary 11.5 implies that $\mathsf{K}$ has also a Beth companion of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Let $f$ be the operation of taking inverses in monoids. Then $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ by Theorem 3.7. Proposition 14.6 yields a unary $g\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ that interpolates $f$ in $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$. By Proposition 14.4 there exist $l,r\in\mathbb{N}$ such that $a^{l}g^{\bm{A}}(a)=a^{r}$ for all $\bm{A}\in\mathsf{K}$ and $a\in\mathsf{dom}(g^{\bm{A}})$. Since $\mathsf{K}$ is a variety and $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ a pp expansion of $\mathsf{K}$, Proposition 10.2 implies that $\mathsf{K}=\mathbb{S}((\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])){\upharpoonright}_{{L}_{\mathsf{K}}})$. Hence, $\mathsf{K}\subseteq\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}})$. Therefore, there exists an extension $\bm{B}$ of the multiplicative monoid $\mathbb{Q}$ of the rationals such that $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$. Then $g^{\bm{B}}$ is total because $g\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})=\mathcal{F}$. Since $g$ interpolates $f$ in $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}$, we obtain $g^{\bm{B}}(2)=f^{\bm{B}}(2)=f^{\mathbb{Q}}(2)=2^{-1}$. Therefore, $2^{l-1}=2^{l}g^{\bm{B}}(2)=2^{r}$. So, $2^{l-1}=2^{r}$ holds in $\mathbb{Q}$. In turn, this implies $l=r+1$. Consequently,
@@
Let $\bm{C}$ and $c\in C$ be as in Claim 14.7. Since $g\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$, there exists $\bm{D}\in\mathsf{K}$ such that $\bm{C}\leqslant\bm{D}$ and $g^{\bm{D}}$ is total. From $c\in\mathsf{dom}(g^{\bm{D}})$ and (68) it follows $c^{r+1}g^{\bm{D}}(c)=c^{r}$. Thus, using $c^{r+1}=c^{r+2}$, we deduce